### v3

# Understanding Mobile App Reviews to Guide Misuse Audits

Abstract
Problem: We address the challenge in responsible computing where an exploitable mobile app is misused by one app user (an abuser) against another user or bystander (victim). We introduce the idea of a misuse audit of apps as a way of determining if they are exploitable without access to their implementation. Method: We leverage app reviews to identify exploitable apps and their functionalities that enable misuse. First, we build a computational model to identify alarming reviews (which report misuse). Second, using the model, we identify exploitable apps and their functionalities. Third, we validate them through manual inspection of reviews. Findings: Stories by abusers and victims mostly focus on past misuses, whereas stories by third parties mostly identify stories indicating the potential for misuse. Surprisingly, positive reviews by abusers, which exhibit language with high dominance, also reveal misuses. In total, we confirmed 156 exploitable apps facilitating the misuse. Based on our qualitative analysis, we found exploitable apps exhibiting four types of exploitable functionalities. Implications: Our method can help identify exploitable apps and their functionalities, facilitating misuse audits of a large pool of apps.
I Introduction
Traditional audits of mobile apps conduct a review of their source code [14, 52]. However, interpersonal misuse arising from app users (instead of app developers) goes unnoticed by such processes. We introduce misuse audit, a process for auditing mobile apps through the reported misuse cases.
A misuse occurs when an app user (abuser) exploits the app to access information of other users or third parties (victims). In particular, a misuse audit can identify cases when the victim (1) doesn’t know about the information access (spying) or (2) may know about the access but is uncomfortable with it. The latter case (often missed by traditional audits) includes incidents of forced consent or when public information (such as on dating apps) is accessed beyond the victim’s level of comfort. We use the term exploitable behavior to denote these two types of information access and use exploitable apps to refer to apps that enable exploitable behavior.
Research [6, 13, 19] shows that exploitable apps may cause discomfort, fear, and potential harm to the victim. Possible ways to prevent this risk include highlighting these apps and their exploitable functionalities to users, app distribution platforms, and app developers. However, identifying these apps is nontrivial since they have a legitimate purpose but are misused by abusers. We propose an approach, MissAuditor, for misuse audits that identify exploitable apps and uncover their exploitable functionalities.
We find that the reviews of an app often describe exploitable functionalities, its misuse (potential and actual), and users’ expectations. Such reviews are evidence of exploitable behavior and can guide audits of such apps. Moreover, app reviews are more valuable than app metadata (such as app descriptions) because metadata indicates only their legitimate purpose and not the actual misuse. In particular, we propose the following research questions:
RQinfo.
What information regarding misuse audit is contained in reviews?
RQidentify.
How can we conduct a misuse audit through app reviews?
RQfunctionality.
What exploitable functionalities are present in audited apps?
Example I shows three reviews (edited for grammar), taken from Apple’s App Store [37] that are relevant to exploitable behaviors. Although our study is based on Apple App Store’s reviews, MissAuditor can be applied to reviews from other sources, including Google’s Play Store [47].
[htb] Example 1: Cases relevant to exploitable behavior Fly on the wall!
(for the AirBeam Video Surveillance app [36])
“with this app, i can spy on my family without them knowing it! it’s such an awesome app!”
This app basically ruined my family to an extent
(for the Life360 app [41])
“My mother made everyone in the family get this app. She freaks out when the app doesn’t do its job because of random obstacles that mess with the location accuracy. Drains the battery and makes my parents paranoid to know where I am at all times. I don’t even do any bad stuff, yet years of trust building are being swept away by the ability to spy on the children of a household. If you’re a parent I highly recommend you don’t get this app because it is extremely uncomfortable to have and it makes parents trust their children less.”
Honest
(for the 3Fun: Threesome & Swingers app [45])
“…A lot of the local people I’ve talked to (Male half of a couple) have been guys who are saying they’re part of a couple, and in all reality are single guys just looking to collect pictures. There is no way to report that that is why you are reporting them. It’s just a boilerplate report feature. I feel there should be a way for the 3Fun community to point out people for bad behavior like this.”
In Example I, the first review, for AirBeam Video [36], addresses the scenario where the app assists a user in accessing a victim’s information without the victim’s knowledge. AirBeam Video is a surveillance app to be installed on the abuser’s device. Hence, the victim may not be an app user but a bystander. The second review, for Life360 [41], complains about the problem of inappropriate access to the user’s location by the user’s mother. Due to the unequal power dynamics between the victim (reviewer in this case) and the abuser (mother in this case), the victim is forced to install apps that violate their privacy. The third review, from 3Fun [45], describes a story of improper access to profile pictures. Even though the profile pictures are public, the victim is uncomfortable with the access. It is common for users to upload such information (pictures in this case) on an app with expectations of how other users would access it. As shown in these three cases, information access may lead to discomfort, fear, or potential harm [13, 19]. Thus, such cases of information access indicate misuse.
I-A Key Findings
We find that reviews contain rich information about apps, which is crucial in auditing them for their misuse potential. For example, not only negative but positive reviews also reveal an app’s exploitable functionalities. Relevant reviews are written by abusers, victims, and third person, each type linguistically distinct, which makes mining reviews challenging (Section II). Moreover, while reporting exploitability, reviews have varying degrees of convincingness and severity, which can be leveraged in misuse audit (Section III). We found that exploitable apps exhibit a variety of exploitable functionalities, ranging from tracking location to monitoring phone activities such as accessing chats, phone contacts, call history, and so on (Section IV).
I-B Contributions and Novelty
Our work’s novelty lies in introducing the problem of misuse audits. We leverage app reviews, which are a large unexplored resource for misuse audits. We contribute to responsible computing in the following manner.
First, we show that app reviews are a viable source of information on its exploitable functionalities. Second, MissAuditor, an approach based on app reviews for misuse audits of apps. Third, a list of exploitable apps along with their alarming reviews revealing exploitable functionalities.
I-C Organization
The rest of this paper is organized as follows. Section II describes our preliminary investigation that shows that app reviews contain evidence useful for audits. Section III describes how to identify exploitable apps using our misuse audit approach, MissAuditor. Section IV shows the procedure to uncover exploitable functionalities of audited apps. Section V lists related work on app reviews and mobile apps. Section VI concludes the paper and Section VII discusses the reproducibility of our findings.
II RQinfo: App Reviews for Misuse Audits
We describe how we collected and investigated reviews for misuse audit, which led to several qualitative findings.
II-A Seed Dataset
Chatterjee et al. [6] identified 2,707 iOS apps as candidates for Intimate Partner Surveillance (IPS), i.e., apps used by a person to spy on their intimate partner. IPS apps are specific to intimate partners and, hence, a subclass of our concept of exploitable apps. We started our analysis from Chatterjee et al.’s IPS candidate list.
During our data collection, 1,687 of these 2,707 apps received at least one review on the Apple App Store, leading to a seed dataset containing 11.57 million reviews. Out of 1,687 apps in the seed dataset, only 210 were confirmed IPS by Chatterjee et al..
II-B Investigating Reviews
Since the seed dataset contains 11.57 million reviews, it is impractical to manually check each review. To find the reviews revealing exploitable behavior, we sampled them using three keywords (spy, stalk, and stealth). From the 5,287 reviews that contain at least one of our keywords, we randomly sampled 1,000 reviews for manual scrutiny. We obtained a simple random sample to investigate if arbitrary reviews reveal any misuse cases. Moreover, we analyzed what information regarding misuse they contain. This sample is diverse: it involves 179 apps with between 1 and 237 reviews each.
Out of 1,000 reviews, we found 403 reviews reporting exploitable behavior and 597 others. We found that reviews reporting exploitable behavior show variation across two dimensions: story and reviewer. Based on the story, we found reviews of the following two types:
exploitable act:
Reviews describe someone performing an exploitable behavior. In such reviews, the reviewer is sure about the app’s exploitable functionality. For example,
“This app is useless and it just helps overprotected parents spy on their sad kids”
potential:
Reviews express the possibility of exploitable behavior. The reviewer may not be sure of the exploitable functionality but identifies risks with the app that can be exploitable in the future. For example,
“…Hate to be a hater. May work well to spy on the kids by ‘accidentally’ leaving iPhone in secret place.”
The functionalities actually or could be misused by abusers (in exploitable act or potential) are called exploitable functionalities. The above two reviews are negative. However, we also found positive reviews indicating an exploitable app. Example II-B shows positive reviews by abusers, who brag about the exploitable functionalities (history tracking in this case) and sometimes express their delight in the misuse. Other reviews are written by victims, who state their concerns and grievances (frustration at the loss of privacy), and some by third persons reporting the misuse of the app.
[htb] Example 2: Types of reviewers reporting exploitable behavior Abuser
“So much better than ¡other app¿. I love the history feature. My kids say it’s creepy and I’m being a stalker. I don’t disagree but it is really nice and convenient to be able to keep track of my kids.”
Victim
“I hate this app so much! My mother is always questioning me and if I delete it she will ground me …No one wants their parents to stalk them!!”
Third Person
“…I don’t feel like parents should track their kids AT ALL. everyone needs a little something called trust and if you don’t have it then your kids will act out and have to become sneaky …I do have this app but only with my friends and we don’t stalk each other …”
Our linguistic analysis reveals that reviews by abusers not only are positive (higher valence than the other two categories) but also illustrate the abuser’s dominance over other parties. Figure 1 shows this difference in the valence and dominance scores between all three reviewers. In general, these relevant reviews show linguistic variations, complicating the task.
Figure 2 shows the relative distribution of 403 relevant reviews across the type of story and reviewer. The third person mostly writes potential stories (73%), whereas abusers and victims mainly describe exploitable acts ($\sim$99and 100%, respectively).
Figure 1: Affective analysis showing valence, arousal, and dominance scores for each type of reviewer. Abusers’ language shows higher dominance and valence than the language used by victims and third persons. Mining app reviews is challenging due to such linguistic variations.
Figure 2: Out of 403 stories manually identified as relevant, abusers and victims mostly write stories indicating exploitable acts ($\sim$99and 100% respectively), whereas third persons write stories indicating potential cases ($\sim$73%).
The above examples indicate that reviews contain rich information about an app’s exploitable behavior. We found two types of stories that are relevant for our purposes. Moreover, such stories are written by three types of reviewers, whose language varies, making the problem of mining these reviews nontrivial. No new types of stories and reviewers were found during our extensive verification of reviews (discussed in Section VI). This indicates that our random sample (1,000 reviews used for this analysis) was a representative set.
III RQidentify: Identifying Exploitable Apps through Misuse Audit
We propose MissAuditor, a review-based approach for misuse audits. Figure 3 shows an overview of the MissAuditor approach. First, we collect reviews from the Apple App Store, as shown for the seed dataset in Section II-A. Second, we label a subset of the collected reviews for alarmingness (defined below) and train a computational model (Section III-A). Third, we apply our model to all reviews to identify exploitable apps (Section III-B). Fourth, we manually examine reviews of some of the identified exploitable apps to find their exploitable functionalities (Section IV).
We envision MissAuditor to be incrementally updated by adding reviews of newly found exploitable apps. Moreover, apps with no reviews can be audited as soon as their reviews arrive.
Figure 3: Overview of MissAuditor.
III-A Alarmingness of Reviews
We find that reviews exhibit two characteristics: (i) convincingness their claims about misuse and (ii) severity or the effect of the misuse. We describe each of these characteristics below.
Convincingness depends on the review’s claim about the app’s exploitability. Some reviews report detailed exploitable behavior and, therefore, are convincing, whereas others are merely suspicions. In Example III-A, the first review is unrelated to exploitable behavior and, hence, is not convincing. The second review describes the reviewer’s suspicion of the app, which may or may not be true (slightly convincing). The third review (by an abuser) confirms the exploitable behavior but lacks details of the exploitable functionalities or victims. In contrast, the other reviews in Example III-A are convincing because they confirm exploitable behavior and mention the location feature, how to set up devices, or the victims being stalked. Extremely convincing reviews also include cases when the app is used for positive purposes, such as tracking family members or pets for safety, but shows the potential to be misused in the future. The reviews that are slightly, moderately, or extremely convincing are relevant to misuse audits. Assigning a convincingness score helps rank all reviews according to the strength of their claims.
Severity measures the effect of exploitable behavior on the victim. Example III-A shows the range of reviews varying in severity. The first review is unrelated to exploitable behavior. Thus, it is not severe. The second review shows that the exploitable act is performed with consent, making this review a slightly severe case. The third review is written by the abuser and lacks the victim’s perspective to analyze the exploitable effect. We assume such acts are performed without consent and consider them moderately severe. The fourth review describes the victim’s misery. The victim even says, “This app has truthfully ruined my teenage years” in the review, which gives solid evidence to be an extremely severe case. Moreover, in the fifth review, the victim complains that others can see when he was last active (also known as last-seen information). This is public information on each profile, but still, the victim is uncomfortable with the access. App developers should be aware of such misuse.
[htb] Example 3: Varying degree of convincingness 1: Not convincing
“It is such a great game, love it so much!”
2: Slightly convincing
“Setup was a breeze. Quicktime 7 pro found it easily. Unfortunately, resolution seems much, much, lower than hoped. Video size can not be adjusted live. Hate to be a hater. May work well to spy on the kids by ‘accidentally’ leaving iPhone in secret place.”
3: Moderately convincing
“This app is perfect for stalking people…”
4: Extremely convincing
This app is awesome for our family to keep track of where everyone is at all times! (You can turn the location off too in case you want to be in stealth mode when buying Christmas presents too.) …Even our dog knows that the alert sound when a family member arrives home means …”
“…I use it to spy on my dogs while I’m at work; so I use it for fun, nothing fancy. My iPad is my camera, and my iPhone is my viewer. …”
bro this app is high key creepy. when i’m with my dad on his days my mom even mentions how she knew everything i was doing and it even made my dad creeped out. if you need this app then ngl yo wack. i don’t want my mom stalking me.”
[htb] Example 4: Varying degree of severity 1: Not severe
“Love the graphics so far it is a great game”
2: Slightly severe
“I love this app, just great because you can time your day accordingly, I like my girlfriend knowing where I am and I love stalking her, we have fun with it.…”
3: Moderately severe
“This app is perfect for stalking people…”
4: Extremely severe
“honestly if you want your kid to rebel against you even more, this is the app for you! This app has truthfully ruined my teenage years all because my mother now has a way of tracking me down 24/7. I couldn’t do the normal teenage things because I was being stalked all day …”
“…i want to share my last seen just to my family and my girlfriend not others. please add new feature in privacy that i can share my last seen to no body except my family and girl friend thanks soo much !”
To capture the above two characteristics, we define the alarmingness of a review as the geometric mean of its convincingness and severity. An app can receive a large number of reviews. Unlike binary classification of reviews, the alarmingness score not only identifies relevant reviews for audit but also ranks them based on the likelihood and danger of the misuse. For an app, the higher the alarmingness of the review, the more useful it is in auditing and identifying misuse.
We created a training set of reviews as follows. From the seed dataset, we randomly selected two sets of about 1,000 reviews each: those that match at least one of our keywords and those that do not. After removing duplicate reviews, we were left with 952 and 932 reviews, respectively, combined into our training corpus of 1,884 reviews. Including both matching and nonmatching reviews in about equal numbers makes our training corpus unbiased toward the keywords.
From these 1,884 reviews, we excluded the reviewers’ identifiers, such as usernames. The task was to rate these reviews for convincingness and severity on a four-point Likert scale (1: not, 2: slightly, 3: moderately, 4: extremely). Two authors were annotators. For reviews rated by both, we computed the average convincingness and severity scores. This annotation study possessed minimal risk and was exempted by the Institutional Review Board (IRB) of our university. We used these annotated reviews as our training data.
We extracted linguistic features of 1,884 reviews through the Universal Sentence Encoder (USE) [5], a widely used approach that has proved effective for app reviews [16, 9]. Using USE features, we trained various multitarget (the targets here being convincingness and severity scores) regression models [4] and found that the Support Vector Regressor (SVR) outperforms others. We considered using the reviews’ metadata, such as ratings and titles, but found the metadata to be unhelpful for identifying misuse and left it out. Not only negative ratings and titles but also reviews with positive ratings and titles indicate misuse (partly discussed in Section II-B). For example, an abuser writes a review with the title “Great app” and provides five-star ratings as the app facilitates misuse. As these nondiscriminatory attributes may confuse the model, we decided not to include them while training. We used only the text of reviews for training.
We leveraged trained SVR to predict convincingness and severity scores of all 11.57 million reviews in the seed dataset. The alarmingness of each review is calculated by taking the geometric mean of its predicted convincingness and severity.
III-B Identifying Exploitable Apps
Using statistical methods, we aggregated the alarmingness of rexviews and ranked all apps according to their reported misuse. From the seed dataset, our model predicted a total of 100 exploitable apps (including false positives). Moreover, our approach is not dependent on the choice of candidate apps and could be applied to any set of apps. To audit additional apps, we applied our model on (i) a dataset of similar apps and (ii) a dataset of 100 popular apps in the utility category. We found not only IPS but also many general-purpose exploitable apps exhibiting multiple exploitable functionalities, which are described below.
III-B1 Similar Apps
For each app determined to be exploitable from the seed dataset, we obtained recommendations for similar apps from the Apple App Store’s “You May Also Like” feature. Through this process, we obtained 788 similar apps. Our motivation in using Apple-recommended apps is that these apps should offer functionalities similar to the apps classified as exploitable. Further, we collected reviews (over the period August 2008 to August 2022) of these 788 apps—we term this the snowball dataset.
Our model predicts 90 apps as exploitable (including false positives) from the snowball dataset. Our evaluation using the snowball dataset reveals that the model yields a recall of 71.60%, which is much higher than the other baseline approaches. In this scenario, a false negative can lead to harm through misuse, whereas a false positive just wastes effort in an unnecessary audit. Hence, high recall is more valuable than high precision.
III-B2 100 Popular Utility Apps
Surveillance apps that can be misused for spying fall under the “Utilities” category, making it an important category to audit. We considered 100 popular utility apps (mentioned on the Apple’s App Store page [46]) and collected their reviews. Since a popular app can have a lot of reviews (over the years), collecting them can be computationally expensive.
From these 100 apps, our approach, MissAuditor, classifies only one app as exploitable. Upon examining its reviews, we found that app to be nonexploitable. We also scrutinized some apps classified as nonexploitable and found their predictions true (based on manual verification of reviews).
Some app categories, such as payment apps, calculators, and so on, are unlikely to be misused. Hence, they don’t form good candidates for identification. Iteratively auditing apps (through MissAuditor) similar to the already identified exploitable apps (as shown in Section III-B1) is a feasible solution to uncover a large exploitable landscape.
III-B3 Relevance of Findings
Some reviews in our datasets are old. For example, the seed dataset was collected over the period of July 2008 to January 2020. To check if our findings are still relevant, we randomly sampled 50 confirmed exploitable apps from the union of the seed and the snowball datasets. We collected new reviews (from January 2020 to April 2024) for these apps and applied our MissAuditor approach to these reviews. Then, we scrutinized their descriptions (to know their basic functionalities) and the alarming reviews (to know misuse cases if any), following the same process as before. We found that the new reviews of three of these 50 apps don’t report misuse. On the other hand, the other 47 apps still possess exploitable functionalities causing potential or actual misuse. A high success rate (94%; 47 of 50) on a random sample implies that most of the identified apps still facilitate misuse.
IV RQfunctionality: Uncovering exploitable Functionalities
We illustrate the idea of uncovering exploitable functionalities using the exploitable apps found from the seed dataset. First, we analyzed an app’s description to understand its functionalities. Second, we analyzed the top alarming reviews (discovered by MissAuditor) for identifying existing misuse and exploitable functionalities. Through this process, we discovered the following types of exploitable functionalities.
Monitoring phone activities.
Some apps monitor a victim’s phone activities, such as browsing history and text messages. Such apps are installed on the victim’s device and activities can be monitored on another synced device. For example, SaferKid Text Monitoring App [44] allows synced devices to monitor call history, web history, texts, and so on.
Audio or video surveillance.
Some apps enable audio or video surveillance without the victim’s knowledge. These apps listen, view, or record a victim’s voice or actions and some of them need not be installed on the victim’s phone. For example, Find My Kids: Parental control [38] is misused to record private conversations between people without them knowing.
Tracking location.
Some Global Positioning System (GPS) apps enable tracking a victim’s phone, making them uncomfortable with access. For example, Find My iPhone [39] is a legitimate app but can be misused to spy on the location of connected devices.
Profile stalking.
Some apps are misused for stalking user profiles or user-generated content (such as text and images). For example, Kik Messaging and Chat App [40] can be misused for stalking images and victims’ other information on the app.
Table IV shows these four types of exploitable functionalities and some alarming reviews reporting them. Some of these reviews are old (2014 or 2012), but we confirmed that similar concerns are being raised in the recent reviews of the same apps. For example, the Find My iPhone [39] app lets its users see the location of the connected devices. Due to unequal power dynamics between the abuser and the victim (say in a family setting) [6], the victim is sometimes forced to connect to such apps and allow their device to be located. In other words, legitimate apps designed for locating loved ones can be misused for exploitable behavior when the requirements for consent [34] are violated. Hence, they are dual-use [6]. Through a qualitative study, Freed et al. [12] found that many find-my-phone applications are intended for anti-theft and safety purposes but are heavily misused against privacy. Chatterjee et al. [6] categorize such apps as spyware, especially in the case of intimate partner surveillance. Just because an app has some legitimate uses doesn’t justify its possible misuse. Hence, we consider such apps to be exploitable. Relying on app reviews highlights both types of exploitable functionalities (always malicious and possibly with legitimate uses) against which future and current app users should be warned. Moreover, app distribution platforms and developers should mitigate privacy risks.
We illustrate the exploitable behavior of the SaferKid Text Monitoring App [44] by installing it on two devices: a parent’s device (iOS version 14.4.1) and a child’s device (Android version 11.0). Figure 5 shows the exploitable features present in this app. Activities on the child’s device can be monitored on the synced parent’s device. Figure 5(a) shows SaferKid exploitable functionalities such as monitoring text messages, web history, and call history. We verified each of these functionalities. Figure 5(b) shows the screen displaying all chats of the victim. Apps such as SaferKid are advertised as safety apps for children but can be secretly or forcefully installed on another device to monitor the user’s activity. Not only parents but anyone can misuse such apps by installing them on a victim’s phone.
(a) Find My iPhone [39] app enables the abuser to track multiple devices at a time.
(b) After selecting a device, abuser can see the directions to victim’s location.
Figure 4: Exploitable functionalities in Find My iPhone [39]. The names and locations of devices are hidden for anonymity.
(a) The SaferKid [44] app provides multiple ways to monitor victim’s phone activities.
(b) Using the SaferKid [44] app, victim’s chats can be seen on the abuser’s phone.
Figure 5: Exploitable functionalities in SaferKid Text Monitoring [44]. The actual chats and device’s name are hidden for anonymity.
TABLE I: Types of exploitable functionalities.
{tabu}
X[2.1,L] X[2.1,L] X[8.5,L] Exploitable Functionality App Example Alarming Review
Monitoring phone activities SaferKid Text Monitoring App [44] …Tracking things like social media, texts, and search history is just a complete disregard of privacy. You have to have trust in your kids …Apps like these shouldn’t be allowed. IF YOU TRUST YOUR KID, DONT DOWNLOAD. (Date: 2019-12-07)
Audio or video surveillance Find My Kids: Parental control [38] This app proves to have a invasion of privacy. Due to the fact if your kids was at a friends house and talking to his friends parents, this app records what is going on and is a invasion of privacy. If your child left their phone downstairs or anywhere and they are playing it can record private conversation between adult and is a unsafe … (Date: 2019-01-16)
Tracking location Find My iPhone [39] …It’s supposed to be used to recover a lost phone, not to religiously stalk your children.…The fact that a mom actually installed this app onto her son’s phone without his knowledge is flat out wrong. …If you’re constantly monitoring your child 24/7, just imagine what your child will do when they go off to college. … (Date: 2014-02-13)
Profile stalking Kik Messaging and Chat App [40] MAKE THIS SAFE PEOPLE WANT TO USE IT BUT DON’T WANT TO BE STALKED OR ABUSED IN THE APP PROTECT THE APP OR PEOPLE WON’T USE IT MY FRIEND HAD A MAN SEND HER A BAD PICTURE IF YOU KNOW WHAT I MEAN. OVERALL THIS APP IS GREAT!!!!!” (Date: 2015-06-24)
V Related Work
We describe previous works focusing on auditing apps, app reviews, and spying through mobile apps.
V-A Auditing and App Reviews
Most research on software audits analyzes the flow and dependencies in source code [14, 52, 29, 27]. Only a few audit studies consider data such as system traces [51]. However, all these studies fail to identify exploitable apps because they focus on the technical aspects of building or running them. Thus, their analysis doesn’t address cases of forced consent and inappropriate access to public information. Our work leverages app reviews, unexplored by prior studies, for misuse audits.
[…]
V-B Spying through Mobile Apps
Chatterjee et al. [6] apply search queries and manual verification based on information such as app descriptions and permissions to identify Intimate Partner Surveillance apps. However, in general, the actual usage deviates from the legitimate purpose shown in app descriptions. To identify such misuse, we focus on the evidence provided in app reviews.
Roundy et al. [33] identify apps used for phone number spoofing and message bombing, which lie outside our scope of exploitable apps. Conversely, exploitable apps include those that enable stalking public information, which lie outside their scope. Roundy et al. use installation data, to uncover spying apps that are installed on infected devices. However, we focus on evidence of exploitable behavior present in app reviews to uncover exploitable apps. Roundy et al. rely upon Norton’s security app [42] to determine which devices are infected. Thus, their approach would miss apps that a general user can leverage to spy.
All these studies, along with a few others [48, 54, 49], investigate how technology is abused for spying. However, they do not support the broader set of exploitable apps. In particular, they do not consider cases when the victim is uncomfortable with the access (e.g., of public information) even if aware of it, along with cases involving family, not only strangers, as abusers.
Some studies address user privacy in the context of user information shared with software developers [32, 3, 22] . They do not consider cases where an app user can exploit the app to access the information of a victim (another app user or bystander).
VI Discussion
We proposed MissAuditor, an approach to automatically analyze app reviews for misuse audits. Specifically, MissAuditor identifies exploitable apps along with their exploitable functionalities from reviews. Doing so reduces the burden of manually installing all apps to perform a misuse audit. MissAuditor predicts exploitable apps from multiple sources. After verification of reviews and apps’ descriptions, we confirmed 156 apps facilitating the misuse. To confirm their misuse, we manually investigated their reviews in two steps. First, we scrutinized their top 50 alarming reviews to confirm any exploitable functionalities. Second, we scrutinized their reviews containing our keywords. Solid misuse cases (in at least one of these two steps) give credence to the exploitability of these 156 apps (full list in the appendix ordered alphabetically).
Below, we describe our process of reporting exploitable apps, threats to validity, and promising future directions.
VI-A Reporting to Apple App Store and App Developers
We informed the Apple App Store about the potential misuse of all exploitable apps that we confirmed above. We found that 17 of them have already been deleted from the Apple Store, possibly because of raised privacy concerns or other unknown reasons.
Only 90 developers (of apps that we confirmed as exploitable) provided their email or chat support on the Apple App Store. We contacted these developers and informed them about the misuse cases (template in appendix). We heard back from 14 of them. Despite asking specific questions, we received generic replies from them. Six of them acknowledged the misuse or informed that their privacy teams take necessary steps in whatever way possible; Five developers assumed that app users (including abusers) would never misuse their apps (an unrealistic assumption); responses from two developers suggested they were not receptive to our feedback; remaining one developer did not answer back after we asked more specific questions.
VI-B Threats to Validity
We now discuss the threats we identify in our work. The identified threats are of two types: (i) the threats that we mitigate; and (ii) the threats that remain.
VI-B1 Threats Mitigated
We mitigated the following threats to validity. To address a potential bias because of our keywords, we constructed a training set that was balanced between reviews that matched and didn’t match our keywords. Using this training set helped our model learn from the context and not from specific keywords. Second, instead of crowd workers, who might not understand the problem well, two authors of this paper annotated the whole training data. Third, the ground truth (of exploitable apps) could be biased if it was formed only using top alarming reviews. We mitigated this bias by scrutinizing reviews from seed and snowball datasets (explained in appendix) containing our keywords, which can have evidence missed by alarming reviews.
VI-B2 Threats Remaining
Now, we describe the threats that remain in our work. First, we investigated only a few thousand apps, which may not represent all apps on the Apple App Store [37]. The performance of MissAuditor may be reduced while testing it on all apps of the Apple App Store. Second, we targeted apps and their reviews only on Apple’s App Store. Upon deployment on other app stores, the performance of our approach can differ. Third, suppose an app distribution platform does not offer recommendations for similar apps. In that case, finding good candidates for exploitable apps will be a challenge. An alternative is to prioritize applying MissAuditor for the apps flagged by app users (victims). Fourth, some reviews in our datasets are old. Due to frequent app updates, the types of misuse may change over time. Although we confirm that our findings are still relevant in Section III-B3, there is a possibility that a few apps may not be exploitable anymore. Fifth, some findings are based on a random review sample because manually reading all reviews is not feasible. The results may vary in the latter case.
VI-C Limitations and Future Directions
We identify the following limitations of this approach, each of which suggests topics for future research. First, MissAuditor may miss some exploitable apps if they do not have alarming reviews at the time of analysis, possibly because they are new apps. However, this limitation can be overcome if MissAuditor is used along with app descriptions. Users could be warned against apps whose descriptions are similar to already identified exploitable apps (or against similar apps found through a recommendation system). Further, MissAuditor can flag these new apps as soon as their alarming reviews arrive. In this way, our computational model can be updated by including newly found exploitable cases in the training set and iteratively identifying a large landscape of misuse.
Second, uncovering exploitable functionalities involves the manual effort of inspecting top alarming reviews. A future direction is to automate this process. Moreover, each functionality is analyzed in isolation from other functionalities. Future audit studies can work on the interdependence of the functionalities facilitating misuse.
Third, app reviews describe perceptions of an app and its misuse, which can be affected by their personal preferences, e.g., how much information access is too much to share. Moreover, cases involving family power dynamics (parent-child and intimate partners), especially regarding forced consent, are nontrivial to analyze user interactions and perceptions. Manual inspection of such reviews by developers can mitigate but not eliminate this limitation. Future studies can explore recommending apps based on a user’s personal preferences.
Fourth, some reviews may falsely claim the exploitable behavior of the app. This includes reviews written by the app’s competitors or by people assuming themselves victims. Identifying such misleading cases is challenging and out of the scope of our study.
Fifth, app reviews can be ambiguous in differentiating between inherently malicious and benign but exploitable functionalities. For example, the Safer Kid Text Monitoring App [44] is designed to keep loved ones safe (benign app) but is misused according to reviews. Future research can distinguish between these benign apps and other malicious ones.
VII Reproducibility
We have released our entire training data along with the computational model on Zenodo [15]. Our appendix provides details of (1) dataset collection and annotation, (2) training and evaluation, and (3) linguistic analysis.
Acknowledgement
Thanks to the anonymous reviewers for their helpful comments. Thanks to the Department of Defense for partially supporting this research under the Science of Security Lablet. NA acknowledges partial support from EPSRC (EP/W025361/1).